# Transform and Entropy Coding in AV2

### v1

@@ III-K Adaptive High Range Coding via Truncated Rice
The coefficients with magnitudes exceeding the thresholds handled by the BR and the LR coding passes are encoded in the High Range (HR) coding pass with bypass bins instead of context-coded bins.
The HR pass uses an adaptive Truncated Rice (TR) code to efficiently represent large values while still maintaining a compact representation for more common, smaller level values. TR code is a combination of Golomb-Rice and Exp-Golomb coding. The idea is to use Golomb-Rice for codes with a shorter unary prefix and switch to Exp-Golomb once the codeword gets too long. By using this combination, the algorithm can take advantage of the flexibility that Golomb-Rice coding has for choosing the Rice Parameter $m$ while keeping the maximum codeword length within 32 bits.
The key feature of this design is the adaptivity of the Rice parameter $m$, which is selected based on a local context value ctxctx that reflects recent statistics of decoded HR coefficient increments. The context is updated incrementally using a moving average of previous HR values $\ell_{\text{HR}}$:
$ctx_{t+1}=(ctx_{t}+\ell_{\text{HR}})\gg 1$ (21)
The moving average captures the relevant trend of the HR level values in a TB in a given scan-order without the need for storing those levels in an intermediate buffer. The context value is initialized to zero at the start of the TB and evolves as HR levels are decoded. The context ctxctx is then mapped to a Rice parameter $m$ using a fixed threshold table of simple powers of two:
[…]
Given $m$, the HR increment level $\ell_{\text{HR}}$ is decoded using a TR code with parameters $(m,k=m+1)$ capped to a unary prefix of length $c_{\text{max}}=\min(m+4,6)$.
The TR code operates as follows:
  • A unary prefix $q$ is read first.
  • If $q<c_{\text{max}}$, the remainder is read using a fixed-length literal of $m$ bits.
  • Else, if $q=c_{\text{max}}$, decode an Exp-Golomb code of order $k$ to represent the residual portion, allowing efficient encoding of arbitrarily large values.
The decoded HR increment is then shifted and added back to the base level to recover the full coefficient value:
$\ell=\ell_{\text{BR/LR}}+(\ell_{\text{HR}}\ll s),\quad s=\begin{cases}1,&\text{if TCQ enabled}\\ 0,&\text{otherwise}\end{cases}$ (23)
@@ III-L Forward Skip Coding (FSC)
Forward Skip Coding (FSC) is introduced in AV2 to improve the efficiency of coefficient coding for TBs that use the 2D identity transform (IDTX). FSC mode is signaled at the CB level for intra blocks and all TBs under the same CB infer this mode. Unlike the coefficient coding logic coupled with traditional trigonometric transforms such as DCT and ADST, FSC directly passes the residual samples after prediction to the entropy coding step and relies on the following rules:
Figure 13: Context derivation neighbors for level and sign coding in FSC. Only two neighboring coefficients (left and above) are used for level context, while three neighbors are used for sign context. Coefficient parsing follows an up-right diagonal scan order.
@@
  • First Position Signaling: A first position index referred to as the beginning of block (BOB) is signaled for FSC-coded TBs to indicate the position of the first coded coefficient. No new syntax is introduced for coding BOB; instead, BOB syntax follows the same logic as EOB (end-of-block) syntax signaling rules in AV2 by using different entropy contexts. Coefficient coding begins from the BOB location and proceeds in forward scan order until the max coefficient index in a given TB. This allows skipping of zero coefficients at the beginning of TBs therefore improving entropy coding throughput.
  • Coding Passes: Coefficients are coded in 3 passes aligned with the ATC Default region logic detailed in Section III-J1. First a BR pass encodes level values [0, 1, 2, $\geq$2] starting from the BOB until the end-of TB in a forward scan order. If the level value is indicated to be greater than 2, then a LR pass is performed starting from the BOB location in forward scan order. The LR pass uses a 4-symbol syntax element and can cover additional level values: [3, 4, 5, $\geq$5]. If the escape character $\geq$5 is not reached, the level coding terminates. If the LR level value reaches to 5, then a HR coding pass is performed to use the Adaptive Bypass Coding logic in Section III-K. The sign information is also coded in the last coding pass.
@@
  • Simplified Context Modeling: Coefficient context derivation is reduced from 5 neighboring coefficients to just 2 (left and above) for coefficient levels and 3 neighboring samples for coding the sign information as shown in Figure 13. The sign and level information are coded using an up-right diagonal scan. Table VIII provides the context numbers used and the derivation rules for each coding pass. The context derivation for level values (BR and LR passes) follows the previously decoded coefficient samples in the neighborhood with a clip, and a particular offset given the TB size. For sign coding, signs of the neighboring 3 samples are used.
    Figure 14: Unified last position signaling used for BOB (beginning-of-block) and EOB (end-of-block) syntax. BOB uses the same position encoding structure as EOB but with different entropy contexts.
  • TB-Level Syntax Skipping: When FSC is enabled, intra-coded CBs infer the IDTX transform type for all sub-TBs. This avoids redundant signaling of a multi-symbol transform type at the TB level, since the parent CB already indicates the transform mode providing compression gains. For inter blocks, the transform type (IDTX) is still signaled at the TB level, as transform candidates are more diverse and TB-level signaling is better suited for inter-coded residuals.
TABLE VIII: FSC Coding Passes and Context Usage
Coding Pass Level Range / CTX Derivation Contexts
Base Range (BR) $ctxId=\mathrm{CLIP}\left(\text{level}_{\text{left}}+\text{level}_{\text{top}},\ 6\right)+\text{TX\_SIZE}_{\text{offset}}$ 21
@@
Low Range (LR) $ctxId=\mathrm{CLIP}\left(\text{level}_{\text{left}}+\text{level}_{\text{top}},\ 6\right)+\text{TX\_SIZE}_{\text{offset}}$ 21
High Range (HR) adaptive bypass coding [Sec. III-K] 0
Sign (with HR) $\text{s}=\text{sign}_{\text{left}}+\text{sign}_{\text{top}}+\text{sign}_{\text{top-left}}$ $\text{ctxId}(s)=\begin{cases}5+\delta&\text{if }s>2\\ 6+\delta&\text{if }s<-2\\ 1+\delta&\text{if }s>0\\ 2+\delta&\text{if }s<0\\ \text{TX\_SIZE}_{\text{offset}}&\text{if }s=0\end{cases}$ $\delta=\begin{cases}2+\text{TX\_SIZE}_{\text{offset}}&\text{if }\text{level}_{\text{curr}}>3\\ \text{TX\_SIZE}_{\text{offset}}&\text{otherwise}\end{cases}$ 27
@@
Overall, FSC serves as an effective and low-complexity enhancement to AV2’s transform and entropy coding framework, especially benefiting IDTX blocks and screen content coding scenarios. It achieves screen-content gains comparable to the Palette tool in AV2, provides coding gains for natural content, and improves performance in near-lossless and lossless coding cases.
@@ IV Coding Gain Estimates
Compression efficiency is measured using the Bjøntegaard Delta (BD) rate for PSNR and VMAF  [bjontegaard2001calculation] under the AOM Common Test Conditions (CTC).
@@ IV-B Lossy Coding Gain Estimates
TABLE XII: BD-rate results (negative is better) for PSNR and VMAF across AI, RA, and LD configurations for each tool at time of adoption into the standard. The values represent estimated individual tool contributions and do not explicitly account for inter-tool dependencies or subsequent tuning. *CoreTX includes gains from core transform partitioning and data-driven transform (DDT) kernel improvements.
@@ IV-C Lossless Coding Gain Estimates
TABLE XIII: Coding gain results (negative is better) across AI, RA, and LD configurations for each lossless tool at time of adoption into the standard.
@@ Appendix A Throughput and CDF Memory Impact
Entropy coding context (CDF) memory usage was analyzed using a standardized entropy coding memory analyzer [AVM-reference] over the course of development. The initial AV2 design increased active CDF storage by 2 KBs. Subsequent simplification reduced this overhead, bringing both RAM and ROM requirements approximately similar to AV1 levels within  0.5 KBs.

### v2

@@ III-K Adaptive High Range Coding via Truncated Rice
Coefficients whose magnitudes exceed the BR and LR ranges are coded in the high-range (HR) stage using bypass bins. AV2 replaces AV1 exponential-Golomb HR coding with an adaptive truncated Rice (TR) code to better match large-magnitude distributions while bounding the worst-case codeword length.
TR follows Golomb-Rice coding when the unary prefix is short and switches to Exp-Golomb when the prefix becomes long. This preserves adaptive Rice-parameter selection while keeping the maximum codeword length within 32 bits. The Rice parameter $m$ is selected from a local context value ctxctx that tracks recently decoded HR increments:
$ctx_{t+1}=(ctx_{t}+\ell_{\text{HR}})\gg 1,$ (21)
initialized to zero per TB. The mapping from ctxctx to $m$ can be written compactly as
$m=\min\!\left(6,\,\max\!\left(1,\,\left\lfloor\log_{2}\!\big(\max(ctx,1)\big)\right\rfloor\right)\right),$ (22)
where $\lfloor\cdot\rfloor$ denotes the floor operator. Given $m$, TR codes the quotient using a unary prefix capped by
$c_{\max}=\min(m+4,6).$ (23)
Let $q$ denote the decoded unary prefix length. If $q<c_{\max}$, TR reads an $m$-bit remainder (Golomb–Rice). If $q=c_{\max}$, TR switches to Exp-Golomb coding of order $k=m+1$ for the remaining value.
The decoded HR increment reconstructs the magnitude as
$\ell=\ell_{\text{BR/LR}}+(\ell_{\text{HR}}\ll s),\quad s=\begin{cases}1,&\text{if TCQ is enabled},\\ 0,&\text{otherwise}.\end{cases}$ (24)
@@ III-L Forward Skip Coding (FSC)
Forward skip coding (FSC) is an AV2 residual-coding mode that improves coefficient-coding efficiency for transform blocks (TBs) coded with the identity transform (IDTX) applied in 2D. FSC is signaled at the coding-block (CB) level, and all TBs within the CB follow FSC coefficient-coding rules:
Figure 14: Context derivation neighborhoods for FSC: two neighbors (left, above) for level contexts and three neighbors (left, above, above-left) for sign contexts.
@@
  • Beginning-of-block (BOB) signaling: A first position index referred to as the beginning of block (BOB) is signaled for FSC-coded TBs to indicate the position of the first coded coefficient. No new syntax is introduced for coding BOB; instead, BOB syntax follows the same logic as EOB (end-of-block) syntax signaling rules in AV2 but uses FSC-specific entropy contexts. Coefficient coding begins from the BOB location and proceeds in forward scan order until the max coefficient index in a given TB, enabling early skipping of leading zeros and improving entropy throughput.
  • Coding passes: Coefficients are parsed in forward scan order starting from the BOB position, and magnitudes are coded in three stages aligned with the Default-region syntax in Section III-J1. BR uses a 4-symbol alphabet $0,1,2,\geq 2]$; if BR escapes, LR codes $3,4,5,\geq 5]$; if LR escapes at $\geq 5$, the remaining magnitude is coded in HR using the adaptive bypass method in Section III-K. Sign information is coded in the final stage together with HR.
@@
  • Simplified context modeling: Coefficient context derivation is reduced from 5 neighboring coefficients to just 2 (left and above) for coefficient levels, and 3 neighboring samples (left, above, above-left) for coding the sign information as shown in Fig. 14. The sign and level information are coded using an up-right diagonal scan. Table VIII provides the context numbers and derivation rules for each coding pass: the context derivation for level values (BR and LR passes) follows the previously decoded coefficient samples in the neighborhood with a clip, and a particular offset given the TB size.
  • TB-level syntax skipping: For intra blocks coded with FSC, the transform type of all sub-TBs is inferred as IDTX, avoiding TB-level transform-type signaling. For inter-coded blocks, IDTX is explicitly signaled at the TB level, and the signaled IDTX implies FSC.
TABLE VIII: FSC coding passes and context usage.
Pass Level range / context derivation Contexts
@@
BR & LR $\mathrm{ctxId}=\mathrm{CLIP}\!\bigl(\mathrm{level}_{\mathrm{left}}+\mathrm{level}_{\mathrm{top}},\,6\bigr)+\mathrm{TX\_SIZE}_{\mathrm{offset}}$ 21
HR adaptive bypass coding (Section III-K) 0
[…]
@@
Overall, FSC provides a low-complexity enhancement for IDTX-coded TBs, with particular benefit for screen and natural content and for near-lossless/lossless operation. It achieves screen-content gains comparable to the Palette tool in AV2, provides coding gains for natural content, and improves performance in near-lossless and lossless coding cases.
@@ IV Coding Gain Estimates
Compression efficiency is measured using the Bjøntegaard Delta (BD) rate for PSNR and VMAF  [4] under the AOM Common Test Conditions (CTC).
@@ IV-B Lossy Coding Gain Estimates
TABLE XII: BD-rate results (negative is better) for PSNR and VMAF across AI, RA, and LD configurations for each tool at time of adoption into AV2. The values represent estimated individual tool contributions and do not explicitly account for inter-tool dependencies or subsequent tuning. *CoreTX includes gains from core transform partitioning and data-driven transform (DDT) kernel improvements.
@@ IV-C Lossless Coding Gain Estimates
TABLE XIII: Coding gain results (negative is better) across AI, RA, and LD configurations for each lossless tool at time of adoption into AV2.
@@ Appendix A Throughput and CDF Memory Impact
Entropy coding context (CDF) memory usage was analyzed using a standardized entropy coding memory analyzer [2] over the course of development. The initial AV2 design increased active CDF storage by 2 KBs. Subsequent simplification reduced this overhead, bringing both RAM and ROM requirements approximately similar to AV1 levels within  0.5 KBs.